OpenStudy (lokiando):

Which product will result in a sum or difference of cubes? a) (x + 7)(x^2 – 7x + 14) b) (x + 8)(x^2 + 8x + 64) c) (x – 9)(x^2 + 9x + 81) d) (x – 10)(x^2 – 10x + 100)

OpenStudy (lokiando):

@UsukiDoll

OpenStudy (usukidoll):

sum of cubes formula \[(a+b)(a^2-ab+b^2)\] difference of cubes formula \[(a-b)(a^2+ab+b^2)\]

OpenStudy (lokiando):

c?

OpenStudy (lokiando):

will give medal

OpenStudy (anonymous):

I think the answer is C.

OpenStudy (lokiando):

you think? mmmmm

OpenStudy (anonymous):

Not really sure on this one. I was thinking either A. or C.

OpenStudy (usukidoll):

ah I see we have to go backwards on this.. so we need all perfect cubes. . no it can't be a.

OpenStudy (anonymous):

Because it can't be factored right? @UsukiDoll

OpenStudy (usukidoll):

8 27 64 \[2^3 = 8 , 3^3 =27, 4^3 = 64, 5^3 = 125\]

OpenStudy (usukidoll):

@200205650 true and also 14 is not a perfect cube

OpenStudy (lokiando):

okay

OpenStudy (lokiando):

so its c?

OpenStudy (usukidoll):

sum of cubes \[(a^3+b^3) = (a+b)(a^2-ab+b^2) \] difference of cubes \[(a^3-b^3) = (a-b)(a^2+ab+b^2) \]

OpenStudy (usukidoll):

these choices are wonky though

OpenStudy (anonymous):

I don't see how you can factor any of them in the answer choices...unless you use the quadratic formula.

OpenStudy (usukidoll):

like .. let a = x let b = 4 \[(a^3-b^3) = (a-b)(a^2+ab+b^2) \] \[(x^3-4^3) = (x-4)(x^2+4x+16) \]

OpenStudy (usukidoll):

I see perfect squares for the last 3 choices that's true.. for the ending part is b^2 . It's like this question isn't following the rules. That's what's getting me mad

OpenStudy (lokiando):

same here

OpenStudy (usukidoll):

let a = x and let b = 9 difference of cubes formula \[(a^3-b^3) = (a-b)(a^2+ab+b^2) \] \[(x^3-9^3) = (x-9)(x^2+9x+81) \]

OpenStudy (anonymous):

From your examples so far @UsukiDoll I would think C. is the right answer.

OpenStudy (usukidoll):

that is messed up... really. 9 isn't a perfect cube to begin with. perfect square yes but not perfect cube

OpenStudy (anonymous):

Ya it isn't...but the way the equations are set up B. and D. don't follow that same pattern.

OpenStudy (usukidoll):

only perfect cubes are allowed for these formulas...>:/ oh I'm missing something XD 9 x 9 x 9 = 729

OpenStudy (usukidoll):

729 is a perfect cube number. that guy is 9

OpenStudy (usukidoll):

ok I got it.. I missed a very important step let a = x and b = 9 \[(x^3-729) = (x-9)(x^2+9x+81)\] \[(x^3-9^3) = (x-9)(x^2+9x+81)\]

OpenStudy (anonymous):

So the answer is C.?

OpenStudy (usukidoll):

that 729 is rewritten as 9^3 in the second line so a = x (the variable) and b = 9 (the number) a^2 = x^2 ab = 9x b^2 = (9)(9) = 81

OpenStudy (usukidoll):

yeah C matches ... on top of that .. the signs match for the difference of cubes. I wasn't working backwards until the very beginning.

OpenStudy (anonymous):

Great job explaining that!

OpenStudy (lokiando):

well thank you guys for your time and patients :)

OpenStudy (usukidoll):

XD patience XDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDD

OpenStudy (usukidoll):

It's past my lunch time too... so that's why my thinking started to go off